OpenStudy (anonymous):

sqrt of 3x+10 = sqrt of x+11 minus 1

OpenStudy (anonymous):

\[\sqrt{3x+10}=\sqrt{x+11}-1\] square both side then we will have 3x+10=x+11-2sqrt(x+11)+1 then we can square it again and we will get a simple equation to solve

OpenStudy (jhannybean):

Haha maybe that's what i did wrong in the other one.

OpenStudy (anonymous):

I don't understand what he said I got the 3x + 10 on the left side but the right side is confusing

OpenStudy (anonymous):

Help meeee @Jhannybean lol

OpenStudy (jhannybean):

i mean (a-b)^2 = a^2-2ab+b^2

OpenStudy (nirmalnema):

9x^2+100+60x=x^2+121+11x-2x-22+1 i think this equation will come after sq..2times

OpenStudy (jack1):

@nirmalnema LHS = correct RHS = not correct

OpenStudy (nirmalnema):

okk..... @Jack1

OpenStudy (jack1):

RHS = x^2 +188 +28x - (48(sqrt(x+11))) - (4x(sqrt(x+11))

OpenStudy (anonymous):

3x+10−−−−−−√=x+11−−−−−√−1

OpenStudy (jack1):

no, should be neg 2(sqrt x+11)

OpenStudy (nirmalnema):

yaa...its -2(sqrt x+11)

OpenStudy (jack1):

sqrt of x+11 minus 1 SQUARED = x + 12 - 2(sqrt(x+11))

OpenStudy (jack1):

its a negative 1 ( a-b )(a -b ) =a^2 -2ab +b^2

OpenStudy (jack1):

so 3x + 10 = x + 12 - 2(sqrt(x+11)) 2x -2 = - 2(sqrt(x+11)) 1-x = sqrt x+11

OpenStudy (jack1):

x^2 - 2x +1 = x+11

OpenStudy (jack1):

x^2 -3x -10 = 0 quadratic rule x = something or x = something else

OpenStudy (jack1):

... -2 and 5...? does that work?

OpenStudy (jack1):

hahahaha! how did we get different answers again?

OpenStudy (anonymous):

the third line of Jhannybean is wrong it should be (x+11) not (x+1)

OpenStudy (nirmalnema):

@Jack1 if u put 5 in the given equation u get 5=3

OpenStudy (jhannybean):

ohh, yeah that's what i meant @johnkwongkwong thanks

OpenStudy (jack1):

so if its 5, sub back into original eqn: sqrt of 3x+10 = sqrt of x+11 minus 1 sqrt 15 = sqrt 16 -1 doesn't work but if x = -2 sub back into original eqn: sqrt of 3x+10 = sqrt of x+11 minus 1 sqrt -4 = sqrt 9 -1 does work so x = -2

OpenStudy (jack1):

@nirmalnema that's the thing hey, x^2 -3x -10 = 0 and there are 2 x values that satisfy this equation but both of those x values won't necessarily satisfy our original equation, so we need to test the answers to find our desired result

OpenStudy (jack1):

@Jhannybean ans = x = -2

OpenStudy (jack1):

no... not tryin to be a tool, sorry

OpenStudy (jhannybean):

im just kidding, :P frustrated that i did it wrong, hahaha.

OpenStudy (jack1):

catch some z's, qt, you'll be runnin circles round everyone again tomorrow

OpenStudy (jhannybean):

lol nah, i need to take a break from OS for a couple days probably...it's so addicting. :|

OpenStudy (jack1):

hmm... maybe socialize with real people a little tomorrow too... ;) good for the soul

OpenStudy (jhannybean):

lol been out all day, actually.

OpenStudy (jack1):

that's ok then slaters

OpenStudy (jhannybean):

Gnight though